Prove ∑ni=11i√≤1n!√∏ni=2(i−1−−−−√)+2∑ni=21i√∑i=1n1i≤1n!∏i=2n(i−1)+2∑i=2n1i using Weierstrass inequality
1
Expert's answer
2019-05-08T08:19:53-0400

We have to prove


"\displaystyle\sum_{i=1}^n{1 \over \sqrt{i}}\leq {1 \over \sqrt{n!}}\displaystyle\prod_{i=2}^n(\sqrt{i-1})+2\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}"

Consider


"\displaystyle\sum_{i=1}^n{1 \over \sqrt{i}}=1+\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}"

Subtract "(2\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}})" from both sides


"\displaystyle\sum_{i=1}^n{1 \over \sqrt{i}}-2\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}=1+\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}-2\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}"

"\displaystyle\sum_{i=1}^n{1 \over \sqrt{i}}-2\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}=1-\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}"

Hence we can rewrite the original inequality as


"1-\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}\leq {1 \over \sqrt{n!}}\displaystyle\prod_{i=2}^n(\sqrt{i-1})"

By the Weierstrass inequality (since "{1 \over \sqrt{i}}\isin (0, 1]" for all "i \in\Z^+") we have


"1-\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}\leq \displaystyle\prod_{i=2}^n(1-{1 \over \sqrt{i}})= \displaystyle\prod_{i=2}^n({\sqrt i-1 \over \sqrt{i}})={1 \over \sqrt{n!}}\displaystyle\prod_{i=2}^n(\sqrt {i}-1)"

For all "i \in\Z^+"


"\sqrt{i}-\sqrt{i-1}={i-(i-1) \over {\sqrt{i}+\sqrt{i-1}}}\leq1"

Then


"\sqrt{i}-1\leq\sqrt{i-1}"

Hence


"1-\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}\leq{1 \over \sqrt{n!}}\displaystyle\prod_{i=2}^n(\sqrt {i}-1)\leq {1 \over \sqrt{n!}}\displaystyle\prod_{i=2}^n(\sqrt {i-1})"

Therefore, we prove that


"\displaystyle\sum_{i=1}^n{1 \over \sqrt{i}}\leq {1 \over \sqrt{n!}}\displaystyle\prod_{i=2}^n(\sqrt{i-1})+2\displaystyle\sum_{i=2}^n{1 \over \sqrt{i}}"